# Nontopological Electromagnetic Hedgehogs

### v1

@@ I Introduction
Theories of self-interacting vector field are, generally, effective field theories; the dynamics of the field [19, 20, 21] and of corresponding solitons may be sensitive to the scale at which the effective theory breaks down and a UV completion is needed [22, 23].
@@
The paper is organised as follows. In Sec. II we introduce the model and derive the field equations and conserved currents. In Sec. III we introduce the field ansatz for solitons and discuss their general properties. We then find the solitons numerically, paying special attention to the nonrelativistic limit and the regime where the thin-wall approximation is applicable. We also touch upon the validity of the effective field theory description for our solutions. In Sec. IV we add the electromagnetic field to the picture of the solitons, and in Sec. V we conclude. Throughout the paper we put $\hbar=c=1$ and use the metric signature (+,,,)(+,-,-,-)( + , - , - , - ).
@@ II Setup
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{2}V^{*}_{\mu\nu}V^{\mu\nu}-\\ -\frac{i\gamma}{2}F_{\mu\nu}W^{\mu\nu}-U(V^{\nu},V^{*\mu})\;,$ (1)
@@
The interaction of massive neutral vector bosons with the gauge field has a clear interpretation in the non-relativistic limit. Namely, the term $\frac{i\gamma}{2}F_{\mu\nu}W^{\mu\nu}$ generates the magnetic dipole moment of a vector particle [25, 26]. As a physical example, the classical field $V_{\mu}$ describes a gas of spin-1 atoms in a Bose-Einstein condensate [27]. In this case, the global $U(1)$ symmetry of the theory (1) corresponds to the $U(1)_{B}$ symmetry for the baryon number conservation. This motivates us to pay special attention to the solitons in the non-relativistic (low-energy) regime of the theory. At the quantum level, eq. 1 is an effective theory that needs a UV completion at a scale $\sim M/\sqrt{\gamma}$ or a scale associated with the vector self-interaction; see, e.g., [28, 23]. Note that if the massive vector field is a dark matter candidate, there is a strong bound on $\gamma$ [29], and the cutoff of the theory (1) is determined by the vector field self-interaction [15].
@@
$\partial_{\mu}F^{\mu\nu}=-i\gamma\partial_{\mu}W^{\mu\nu}\;,$ (2)
which, in the absence of external fields, can be integrated to yield
$F^{\mu\nu}=-i\gamma W^{\mu\nu}\;.$ (3)
@@
where $\tilde{\alpha}=\alpha-\gamma^{2}$ and $\tilde{\beta}=\beta+\gamma^{2}$.
Let us discuss the possibility of enhancing the Lagrangian (1) with the P-odd term $\frac{i\tilde{\gamma}}{2}\tilde{F}_{\mu\nu}W^{\mu\nu}$, where $\tilde{F}_{\mu\nu}=\varepsilon_{\mu\nu\rho\lambda}F^{\rho\lambda}$ and $\tilde{\gamma}$ is a dimensionless coupling constant. In the non-relativistic limit, this interaction provides the vector boson with the electric dipole moment. Considering this coupling instead of $\frac{i\gamma}{2}F_{\mu\nu}W^{\mu\nu}$ leads again to the effective theory (LABEL:eff_model) with $\tilde{\alpha}=\alpha+4\tilde{\gamma}^{2}$ and $\tilde{\beta}=\beta-4\tilde{\gamma}^{2}$. Note that both P-even and P-odd couplings provide the P-even contributions to the effective Lagrangian. This is because $W_{\mu\nu}$ (or $\tilde{W}_{\mu\nu}=\varepsilon_{\mu\nu\rho\sigma}W^{\rho\sigma}$) gets squared when moving from (1) to (LABEL:eff_model). If both $\gamma$ and $\tilde{\gamma}$ are nonzero, the P-odd interaction term $\frac{\gamma\tilde{\gamma}}{2}W^{\mu\nu}\tilde{W}^{\mu\nu}$ should also be added to the vector field Lagrangian. In this paper, we will be interested in configurations supporting the electric (or magnetic) field in the center-of-mass frame, hence the latter term is identically zero for our solitons.
@@ III.1 Condensate
Here w is the angular frequency in units of the field mass, and the sign in the denominator follows the sign of $\tilde{\alpha}$. In the non-relativistic limit, $\mathrm{w}\to 1$, this condensate approaches the classical ground state $\tilde{V}^{\mu}=0$. At $\mathrm{w}<1$, the solution (12) describes weakly-interacting vector particles in a spatially-homogeneous bound state. The existence of such bound state is important for the non-relativistic solitons that are physically formed out of the gas of weakly-interacting particles. Thus, we require ±$\pm 1+\kappa>0$ for our solitons. Another way to obtain this constraint is to study directly the non-relativistic limit of solitons with a particular ansatz; see, e.g., [7].
@@
where $u$ and $v$ are real constants depending on $\mathrm{w}$. Using equations of motion (11), we obtain that the bound-state solution of the form (13) exists at $-1<\kappa<0$. The existence of this bound state is important for the kinematic stability of the solitons. Indeed, the solitons with $E<MQ$ are often associated with the thin-wall limit [3]. Near this limit, the bulk of the soliton, which makes the dominant contribution to its charge and energy, is close to the respective condensate solution. If the latter is kinematically stable, so is the soliton. As we will see, the ansatz (13) describes the bulk of our solitons in the thin-wall regime. Thus, we require $-1<\kappa<0$, which, together with the bound ±$\pm 1+\kappa>0$, also implies that $\tilde{\alpha}>0$.
@@ III.2 Solitons
The rest of the paper is dedicated to solving numerically eq. 15 with the solitonic boundary conditions, and discussing the properties of the obtained solutions. Our procedure is as follows. We fix the value of $\kappa$ from the range $-1<\kappa<0$, in which one expects solutions both in the non-relativistic and the thin-wall (kinematically stable) limits. Using the shooting method, we first find solutions close to the non-relativistic limit, $\mathrm{w}\rightarrow 1$. In this limit, the field amplitudes tend to zero, and the validity of the effective theory is ensured.
@@
Let us discuss the limit $\mathrm{w}\to\mathrm{w}_{min}$ in more detail. There are two types of solutions near this limit; both of them are kinematically stable. Solutions of the first type—the proper thin-wall solitons—are found in the range $-1<\kappa\lesssim 0.6$. For such solitons, the region of the central depression of the fields is followed by a broad bulk where the magnitudes of $u$, $v$ approach the values of the condensate solution (13) with the same frequency. Outside the bulk, the fields rapidly vanish. This structure is illustrated in Fig. 2. The size of the soliton, its total charge and energy grow indefinitely as $\mathrm{w}\to\mathrm{w}_{min}$.
@@
Solutions of the second type are found in the range $-0.6\lesssim\kappa\lesssim 0$. They do not approach the thin-wall limit even though the condensate (13) still exists. The obstruction to the limit lies in the structure of the differential equations (15). To see it, we express $v^{\prime}$ in terms of $u$, $\chi\equiv u^{\prime}$ and $v$ (which brings eq. 15 to the standard form for the Runge-Kutta method):
@@
Having obtained the family of solitons parameterised by w, we can study relations between their charge $Q$ and energy $E$. For example, Fig. 4 shows the parametric dependence of $E/M-Q$ on $Q$, at $\kappa=-0.55$. We see the lower branch heading towards the thin-wall limit, and the upper branch connecting to the non-relativistic limit. The branches are joined at a point where $\partial E/\partial\omega=\partial Q/\partial\omega=0$. The qualitatively same dependencies are obtained for other allowed values of $\kappa$ (for $\kappa\gtrsim-0.4$, however, the soliton develops the cusp at $E/M-Q>0$, and the kinematically stable region is inaccessible). The behavior of $E$ and $Q$ is quite similar to the one in the Friedberg-Lee-Sirlin model of Q-balls made of two scalar fields [30].
@@
Figure 4: $E/M-Q$ as a function of $Q$ for vector solitons (14) in the theory (LABEL:eff_model) with $\tilde{\alpha}=1$ and $\kappa=\tilde{\beta}/|\tilde{\alpha}|=-0.55$.
@@ III.3 Applicability of classical theory
Let us discuss the validity of the solutions obtained above. We work in the regime $1\gtrsim\alpha\gg\gamma$. The first inequality is required by the semiclassical approximation. The second inequality is motivated by considering $V^{\mu}$-particles as dark matter candidate [29]. The strong coupling scale of the theory with the potential (4) can be estimated as $\Lambda\sim M\sqrt{4\pi}/\alpha^{1/4}$ [31]. On the other hand, from eq. 10 it follows that the physical field $V^{\mu}$ scales as $1/\alpha^{1/2}$. Requiring the physical field amplitude stay below the cutoff leads to the bound $\alpha>1/(4\pi)^{2}$. The above numerical analysis shows that kinematically-stable solitons with $|\kappa|$, w sufficienty close to 1 (such as the one in Fig. 1) are within the validity of the effective theory provided that $\alpha$ is not very small.
The charge (16) and energy (17) of these solitons satisfy $Q\gg 1$, $E/M\gg 1$, meaning that the soliton is composed of many quanta of the field $V^{\mu}$. Furthermore, the size of the soliton is much larger than its Compton wavelength $\sim E^{-1}$. Finally, from Fig. 4 we see that $|E-MQ|\ll MQ$, meaning that the binding energy of $V^{\mu}$-particles inside the soliton is small compared to their rest energy, and the particles are non-relativistic.
Note that the solitons with the cusp, such as the one presented in Fig. 3, are beyond the applicability limit of the effective theory, since they have large quantum corrections [19].
@@
According to eq. 3, the vector solitons can confine the electromagnetic field. As discussed above, we choose $\gamma\ll\alpha\lesssim 1$. At $-1<\kappa<0$, we have found solitons which are non-relativistic, kinematically stable, and have a wide range of physical parameters $Q$ and $E$. These spherically-symmetric solutions support radial fields—“hedgehogs”, see Fig. 5 for illustration. For the P-even interaction between $F_{\mu\nu}$ and $W^{\mu\nu}$, it is the radial electric field trapped in the bulk of the soliton, while for the dual theory ($\gamma=0$, $\tilde{\gamma}\neq 0$) it is the radial magnetic field. Note that these nontopological solitons are not magnetic monopoles, since the magnetic field decreases exponentially fast at large distances. This localisation of the field is the feature of the theory (1).
@@ V Conclusion
Another natural generalisation is to use gravity as a binding force of the vector soliton, rather than the field self-interaction. This would allow one to disentangle the parameters of the soliton from its coupling to the electromagnetic field. Next, the theory (LABEL:eff_model) is an effective theory and needs a UV completion by, e.g., a Higgs-like field. Solitons in a resulting theory—“Proca-Higgs balls” and their gravitating counterparts—are a subject of recent research [23, 34]. Finally, one can study small perturbations of the solutions presented here. In particular, it is important to analyse linear classical stability of the vector solitons and of the electromagnetic field trapped by the soliton.

### v2

@@ I Introduction
Theories of self-interacting vector field are effective field theories. In the exceptional case of non-Abelian gauge theory, the well-known weakly coupled UV completion involves the Higgs mechanism, but this approach is not applicable in the general case. [^1] For example, spin-1 atoms and $\rho$-mesons are composite particles in the complete strongly coupled theories. Moreover, as was shown in [28, 8, 9], pathologies in models of self-interacting Proca field may appear already at the classical level. This does not preclude viewing relativistic massive vector theories as low-energy effective theories. Solitons and other classical configurations can be studied in this framework [22, 19], which we adopt in this paper.
@@
The paper is organised as follows. In Sec. II we introduce the model and derive the field equations and conserved currents. In Sec. III we introduce the field ansatz for solitons and discuss their general properties. We then find the solitons numerically, paying special attention to the nonrelativistic limit and the regime where the thin-wall approximation is applicable. We also touch upon the validity of the effective field theory description for our solutions. In Sec. IV we add the electromagnetic field to the picture of the solitons, and in Sec. V we conclude. Throughout the paper we put $\hbar=c=1$ and use the metric signature (+,,,)(+,-,-,-).
@@ II Setup
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{2}V^{*}_{\mu\nu}V^{\mu\nu}+\\ +\frac{i\gamma}{2}F_{\mu\nu}W^{\mu\nu}-U(V^{\nu},V^{*\mu})\;,$ (1)
@@
The real interaction term $\frac{i\gamma}{2}F_{\mu\nu}W^{\mu\nu}$ was first introduced in [12]. In particle physics, this interaction describes three-vector-boson coupling which was considered in Electroweak phenomenology [17] and studied in a low-energy effective theory of strongly interacting spin-1 particles [14]. In the non-relativistic limit the term $\frac{i\gamma}{2}F_{\mu\nu}W^{\mu\nu}$ generates the magnetic dipole moment $(\gamma/2M)$ [30, 24] and the electric quadrupole moment $(-\gamma/M^{2})$ [17, 14] of a vector particle.
As a physical example, the classical field $V_{\mu}$ describes a gas of spin-1 atoms in a Bose-Einstein condensate [13]. In this case, the global $U(1)$ symmetry of the theory (1) corresponds to the $U(1)_{B}$ symmetry for the baryon number conservation. This motivates us to pay special attention to the solitons in the non-relativistic (low-energy) regime of the theory. At the quantum level, the model (1) is an effective theory that is valid below the cutoff scale $\sim M/\sqrt{\gamma}$ or a scale associated with the vector self-interaction. Note that the model can not be considered as a low-energy limit of some renormalizable non-Abelian gauge theory. One possibility for a UV completion is to match the model to some strongly coupled theory at low energies [14]. We will follow the approach of [22, 31, 19] and study classical solutions in the model assuming that the coupling constants in (1) are independent. Note also that if the massive vector field is a dark matter candidate, there is a strong bound on $\gamma$ [7], and the cutoff of the theory is determined by the vector field self-interaction [2].
@@
$\partial_{\mu}F^{\mu\nu}=i\gamma\partial_{\mu}W^{\mu\nu}\;,$ (2)
which, for configurations with trivial boundary conditions at spatial infinity, can be integrated to yield
$F^{\mu\nu}=i\gamma W^{\mu\nu}\;.$ (3)
@@
where $\tilde{\alpha}=\alpha+\gamma^{2}$ and $\tilde{\beta}=\beta-\gamma^{2}$.
Let us discuss the possibility of enhancing the Lagrangian (1) with the P-odd term $\frac{i\tilde{\gamma}}{2}\tilde{F}_{\mu\nu}W^{\mu\nu}$, where $\tilde{F}_{\mu\nu}=\varepsilon_{\mu\nu\rho\lambda}F^{\rho\lambda}$ and $\tilde{\gamma}$ is a dimensionless coupling constant. In the non-relativistic limit, this interaction provides the vector boson with the electric dipole and magnetic quadrupole moment. Considering this coupling instead of $\frac{i\gamma}{2}F_{\mu\nu}W^{\mu\nu}$ leads again to the effective theory (LABEL:eff_model) with $\tilde{\alpha}=\alpha-4\tilde{\gamma}^{2}$ and $\tilde{\beta}=\beta+4\tilde{\gamma}^{2}$. Note that both P-even and P-odd couplings provide the P-even contributions to the effective Lagrangian. This is because $W_{\mu\nu}$ (or $\tilde{W}_{\mu\nu}=\varepsilon_{\mu\nu\rho\sigma}W^{\rho\sigma}$) gets squared when moving from (1) to (LABEL:eff_model). If both $\gamma$ and $\tilde{\gamma}$ are nonzero, the P-odd interaction term $\frac{\gamma\tilde{\gamma}}{2}W^{\mu\nu}\tilde{W}^{\mu\nu}$ should also be added to the vector field Lagrangian. In this paper, we will be interested in configurations supporting the electric (or magnetic) field in the center-of-mass frame, hence the latter term is identically zero for our solitons.
@@ III.1 Condensate
Here w is the angular frequency in units of the field mass, and the sign in the denominator follows the sign of $\tilde{\alpha}$. In the non-relativistic limit, $\mathrm{w}\to 1$, this condensate approaches the classical ground state $\tilde{V}^{\mu}=0$. At $\mathrm{w}<1$, the solution (12) describes weakly-interacting vector particles in a spatially-homogeneous bound state. The existence of such bound state is important for the non-relativistic solitons that are physically formed out of the gas of weakly-interacting particles. Thus, we require $\pm 1+\kappa>0$ for our solitons. Another way to obtain this constraint is to study directly the non-relativistic limit of solitons with a particular ansatz; see, e.g., [26].
@@
where $u$ and $v$ are real constants depending on $\mathrm{w}$. Using equations of motion (11), we obtain that the bound-state solution of the form (13) exists at $1<\kappa<0$. The existence of this bound state is important for the kinematic stability of the solitons. Indeed, the solitons with $E<MQ$ are often associated with the thin-wall limit [11]. Near this limit, the bulk of the soliton, which makes the dominant contribution to its charge and energy, is close to the respective condensate solution. If the latter is kinematically stable, so is the soliton. As we will see, the ansatz (13) describes the bulk of our solitons in the thin-wall regime. Thus, we require $1<\kappa<0$, which, together with the bound $\pm 1+\kappa>0$, also implies that $\tilde{\alpha}>0$.
@@ III.2 Solitons
The rest of the paper is dedicated to solving numerically eq. 15 with the solitonic boundary conditions, and discussing the properties of the obtained solutions. Our procedure is as follows. We fix the value of $\kappa$ from the range $1<\kappa<0$, in which one expects solutions both in the non-relativistic and the thin-wall (kinematically stable) limits. Using the shooting method, we first find solutions close to the non-relativistic limit, $\mathrm{w}\rightarrow 1$. In this limit, the field amplitudes tend to zero, and the validity of the effective theory is ensured.
@@
Let us discuss the limit $\mathrm{w}\to\mathrm{w}_{min}$ in more detail. There are two types of solutions near this limit; both of them are kinematically stable. Solutions of the first type—the proper thin-wall solitons—are found in the range $1<\kappa\lesssim 0.6$. For such solitons, the region of the central depression of the fields is followed by a broad bulk where the magnitudes of $u$, $v$ approach the values of the condensate solution (13) with the same frequency. Outside the bulk, the fields rapidly vanish. This structure is illustrated in Fig. 2. The size of the soliton, its total charge and energy grow indefinitely as $\mathrm{w}\to\mathrm{w}_{min}$.
@@
Solutions of the second type are found in the range $0.6\lesssim\kappa\lesssim 0$. They do not approach the thin-wall limit even though the condensate (13) still exists. The obstruction to the limit lies in the structure of the differential equations (15). To see it, we express $v^{\prime}$ in terms of $u$, $\chi\equiv u^{\prime}$ and $v$ (which brings eq. 15 to the standard form for the Runge-Kutta method):
@@
Having obtained the family of solitons parameterised by w, we can study relations between their charge $Q$ and energy $E$. For example, Fig. 4 shows the parametric dependence of E/MQE/M-Q on $Q$, at $\kappa=-0.55$. We see the lower branch heading towards the thin-wall limit, and the upper branch connecting to the non-relativistic limit. The branches are joined at a point where $\partial E/\partial\omega=\partial Q/\partial\omega=0$. The qualitatively same dependencies are obtained for other allowed values of $\kappa$ (for $\kappa\gtrsim-0.4$, however, the soliton develops the cusp at E/MQ>0E/M-Q>0, and the kinematically stable region is inaccessible). The behavior of $E$ and $Q$ is quite similar to the one in the Friedberg-Lee-Sirlin model of Q-balls made of two scalar fields [15].
@@
Figure 4: E/MQE/M-Q as a function of $Q$ for vector solitons (14) in the theory (LABEL:eff_model) with $\tilde{\alpha}=1$ and $\kappa=\tilde{\beta}/|\tilde{\alpha}|=-0.55$.
@@ III.3 Applicability of classical theory
Let us discuss the validity of the solutions obtained above. We work in the regime $\gamma\ll\alpha\lesssim 1$. One of these inequalities is required by the semiclassical approximation: as discussed above, the classical action of the theory with the potential (4) scales as $|\alpha|^{-1}$, hence quantum corrections are suppressed by the powers of small coupling $\alpha$ and additional loop factor $1/(4\pi)^{2}$. The strong coupling scale of the theory is estimated as $\Lambda_{\alpha}\sim M\sqrt{4\pi}/\alpha^{1/4}$ [32]. It should be larger than the interacting energy of particles in the soliton, which is satisfied automatically for our solutions at small coupling constants. Note that, even though the magnitude of the field in the soliton may exceed the cutoff scale, $V^{\mu}\sim M/\sqrt{\alpha}>\Lambda_{\alpha}$, this does not mean the breakdown of the semiclassical approximation, which follows from the loop expansion controlled by the powers of $\alpha$. [^2] A similar procedure is used to validate the semiclassical analysis of the Higgs mechanism for parametrically large vacuum expectation values [10]. The second inequality is motivated by considering $V_{\mu}$-particles as a dark matter candidate. For example, the direct constraint on the magnetic dipole moment of electromagnetically interacting vector particles with masses of order 100 MeV is $\gamma<10^{-8}$ [7].
The charge (16) and energy (17) of our solutions satisfy $Q\gg 1$, $E/M\gg 1$, meaning that the solitons are composed of many quanta of the field $V^{\mu}$. Furthermore, the size of the soliton is much larger than its Compton wavelength $\sim E^{-1}$. Finally, from Fig. 4 we see that $|E-MQ|\ll MQ$, meaning that the binding energy of $V^{\mu}$-particles inside the solitons is small compared to their rest energy, and the particles are non-relativistic.
Note, finally, that the solitons with the cusp, such as the one presented in Fig. 3, are beyond the applicability limit of the effective theory, since they have large quantum corrections [28].
@@
According to eq. 3, the vector solitons can confine the electromagnetic field. As discussed above, we choose $\gamma\ll\alpha\lesssim 1$. At $1<\kappa<0$, we have found solitons which are non-relativistic, kinematically stable, and have a wide range of physical parameters $Q$ and $E$. These spherically-symmetric solutions support radial fields—“hedgehogs”, see Fig. 5 for illustration. For the P-even interaction between $F_{\mu\nu}$ and $W^{\mu\nu}$, it is the radial electric field trapped in the bulk of the soliton, while for the dual theory ($\gamma=0$, $\tilde{\gamma}\neq 0$) it is the radial magnetic field. Note that these nontopological solitons are not magnetic monopoles, since the magnetic field decreases exponentially fast at large distances. This localisation of the field is the feature of the theory (1).
@@
Since $\mathrm{w}\approx 1$, the solution can be described in the non-relativistic language where it has a clear physical interpretation. In the non-relativistic regime, one can re-write the equation of motion for $V^{i}$ in the form of the Gross-Pitaevsky equation with additional terms describing magnetic dipole and electric quadrupole interactions, and use the transversality condition $\partial_{\mu}V^{\mu}=0$ to exclude $V_{0}$ [30]. Note that the magnetic dipole moment vanishes on the ansatz (14) due to antisymmetry of $\epsilon_{ijk}$. This means that the leading electromagnetic interaction is between the electric field and the electric quadrupole moment of the vector field. It is this interaction that provides the soliton with the structure of the electric hedgehog shown in Fig. 5.
@@ V Conclusion
Another natural generalisation is to use gravity as a binding force of the vector soliton, rather than the field self-interaction. This would allow one to disentangle the parameters of the soliton from its coupling to the electromagnetic field. Next, one can add a Higgs-like field in the theory (LABEL:eff_model) to obtain configurations with lower values of $\mathrm{w}$. Solitons in a resulting theory—“Proca-Higgs balls” and their gravitating counterparts—are a subject of recent research [19, 4]. Finally, one can study small perturbations of the solutions presented here. In particular, it is important to analyse linear classical stability of the vector solitons and of the electromagnetic field trapped by the soliton.